OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -4 + 5 + 14 + 23 + ... + 131

OpenStudy (anonymous):

It seems to go by 9s.

OpenStudy (asnaseer):

it will help if you wrote down the partial sums

OpenStudy (asnaseer):

e.g. sum to 1 term is -4 sum to 2 terms is -4 + 5 = -4 + (-4 + 9) sum to 3 terms is -4 + 5 + 14 = -4 + (-4 + 9) + (-4 + 9 + 9) etc can you spot a pattern?

OpenStudy (anonymous):

no i do not

OpenStudy (asnaseer):

try writing a few more partial sums and see if you can spot how this is growing

OpenStudy (anonymous):

wait i do so the next one would be 32?

OpenStudy (asnaseer):

correct

OpenStudy (asnaseer):

and 32 can also be written as: -4 + 9 + 9 + 9 + 9

OpenStudy (asnaseer):

so each term in this sum can be written as -4 plus some multiple of 9

OpenStudy (anonymous):

ok thank you so much

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

can you help with a couple more things?

OpenStudy (asnaseer):

if you post each question in the list on the left then I am sure there are plenty of people on this site who will come to help you. I will try as well - but I also need to do some moderator duties, so I am not always fully available.

OpenStudy (anonymous):

ok no problem